# A Survey on Data Selection for Language Models

### v2

@@ 10.5 Considerations when applying data selection to your setting
The first step in determining appropriate actions for data selection is to determine whether you are in a setting where you wish to increase the distribution coverage from your data, or if you wish to select a subset from your data. For pretraining, the goal will usually be to increase coverage, unless the model has a well defined downstream use case (e.g. for medical domain), in which case the goal can be to select a subset of general data which is most similar to the downstream use case. For multitask training and instruction tuning, the goal is generally to increase coverage because the goal of these settings is to produce a capable general purpose model. Additionally, depending on the number of parameters in the model you plan to train, you may have a range for the desired number of training tokens according to your compute budget (Kaplan et al., 2020; Hoffmann et al., 2022; Penedo et al., 2023). The various filter sensitivities will need to be adjusted depending on the number of tokens in the raw data in order to achieve the desired number of training tokens. For task-specific fine-tuning, the setting depends on how large the task-specific dataset is. Generally, if the dataset is relatively small, the goal of data selection will be to increase coverage over the expected test distribution. However, if the task-specific dataset is very large then models can plausibly model the distribution with fewer data points by removing redundant information from the dataset, and costs can be reduced.
The next step in the decision making is with regards to how much information you have on the test distribution as this will determine which utility functions available to you. Depending on whether you have data sampled from the test distribution, you may be able to find an appropriate model-based utility function, otherwise you will need to determine appropriate heuristics. Furthermore, the available utility functions may be limited by the amount of compute available.
@@ 11.1 Accelerating research on data selection
Research on data can be a very slow and tedious process, potentially requiring significant investments of time, compute, and human effort. Here we describe four directions of future research that can reduce the effort required, and the barriers to entry, for research in data selection.
@@
One direction that can enable more researchers to participate, as well as allowing for faster iteration is scaling down the current dataset sizes and models. For example, Kaddour (2023) demonstrate that it is possible to train a model on 745x less data, but lose only 1.9% performance on GLUE (Wang et al., 2018a) and 2.5% performance on super-natural instructions (Wang et al., 2022b). Additionally, Polo et al. (2024) show that a wide variety of evaluation benchmarks can be distilled into only 100 data points, while maintaining an accuracy within  2% of the true accuracy. One area that is still uncertain is whether the results from small models scale up to larger models, or if there is any way to predict the results on large models, based on experiments from small models. This is an active area of research.
@@
Preliminary works have been done in this direction, including some works on data attribution and valuation (Ghorbani & Zou, 2019; Ilyas et al., 2022), and data measurements (Mitchell et al., 2022). Works on data valuation and attribution (discussed in §9) attempt to better understand how each data point affects a models predictions, but may be used as inspiration for methods that directly evaluate the data itself. Additionally, there are a number of metrics that directly measure the intrinsic characteristics of data including distance, density, diversity, tendencies (mean, median, mode), and association (correlation, mutual information), but these measures have yet to be connected with a model’s performance on downstream data (Mitchell et al., 2022).
Developing metrics that directly evaluate data can significantly increase the time spent on method development and increase the amount of exploratory work possible. Furthermore, removing the need to train models (which are very costly) can allow for lower-resourced organizations and individuals to contribute to the field, improving diversity and inclusion in the process of large model development.
@@
Another high-impact direction for increasing the rate of research on data selection is the development of data-centric benchmarks and challenges. One hurdle of advancing data selection is that works are done across a wide variety of settings and cannot be directly compared. Specifically, much of data selection research using heuristics filters is performed as part of collecting a new dataset. This makes it impossible to compare selection methods with another dataset as they may contain different quantities of data, and sometimes even use varying model architectures and sizes. In other domains, such as vision and vision+language, challenges and benchmarks have been developed that provide well-founded evaluation settings that allows researchers from many institutions to participate in a combined effort to push the frontier of knowledge and capabilities in data selection (discussed in "Increasing the rate of data selection research" in §8). Creating similar challenges for language models can go a long way in encouraging progress.
@@ 11.3 Shifting compute time from model training to data processing
It is now well known that training data plays a very significant importance in the performance of a model, therefore creating an increased desirability to spend more compute on data processing and selection (and possibly away from model training). To improve beyond what is possible only through heuristics, it may be necessary to utilize more expensive methods.
One example of how this may be accomplished is through the use of methods which have previously only been used in low-data settings (e.g. alignment) and applying them to settings with greater amounts of data (e.g. pretraining). For example, methods inspired by model-based filtering, or reward model-reweighting techniques can be used to improve the “quality” of pretraining data. Additionally, techniques for generating synthetic data can be used to create more diverse training data Bradley et al. (2023) and methods previously used only for red-teaming (Samvelyan et al., 2024) can be used to generate training data for more robust models. As compute gets cheaper, it becomes more feasible to spend the cost of directly applying the existing methods. However, another avenue worth exploring is using cheap approximations for utility functions, such as computing similarity using bag-of-n-grams and hash functions rather than from model representations.
Another method for improving data that would require large amounts of compute is additive data selection. Currently, most data selection methods (especially for pretraining) remove undesirable content from data points. However, it is possible that in the future it will be beneficial to have additive data selection methods. For example, it would be possible to use a language model to infill information within a data point. One example would be to include explanations for some phrases: “e.g. the basketball player wasn’t able to move due to the double dribble rule so he passed the ball to his teammate” where the italicized text has been added by an infilling model.
@@ Appendix A Heuristic Filtering Details
  • If the line is a counter (e.g. “3 likes”)
@@
  • At the beginning of the line (e.g. “sign-in”)
  • At the end of the line (e.g. “Read more…”)
  • Anywhere in the line (e.g. “items in cart”)
@@
  • have an word from the "List of Dirty, Naughty, Obscene, or Otherwise Bad Words" [^35] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words

### v3

@@ 10.5 Considerations when applying data selection to your setting
The first step in determining appropriate actions for data selection is to determine whether you are in a setting where you wish to increase the distribution coverage from your data, or if you wish to select a subset from your data. For pretraining, the goal will usually be to increase coverage, unless the model has a well-defined downstream use case (e.g., for medical domain), in which case the goal can be to select a subset of general data which is most similar to the downstream use case. For multitask training and instruction tuning, the goal is generally to increase coverage in order to ultimately produce a capable general purpose model. Additionally, depending on the number of parameters in the model you plan to train, you may have a range for the desired number of training tokens according to your compute budget (Kaplan et al., 2020; Hoffmann et al., 2022; Penedo et al., 2023). The various filter sensitivities will need to be adjusted depending on the number of tokens in the raw data in order to achieve the desired number of training tokens. For task-specific fine-tuning, the setting depends on how large the task-specific dataset is. Generally, if the dataset is relatively small, the goal of data selection will be to increase coverage over the expected test distribution. However, if the task-specific dataset is very large then models can plausibly model the distribution with fewer data points by removing redundant information from the dataset, and costs can be reduced.
The next step in the decision making depends on how much information you have about the test distribution, as this will determine which utility functions are available to you. Depending on whether you have data sampled from the test distribution, you may be able to find an appropriate model-based utility function, otherwise you will need to determine appropriate heuristics. Furthermore, the available utility functions may be limited by the amount of compute available.
@@ 11.1 Accelerating research on data selection
Research on data can be a very slow and tedious process, potentially requiring significant investments of time, compute, and human effort. Here we describe four directions of future research that can reduce the effort required and the barriers to entry for research in data selection.
@@
One direction that can enable more researchers to participate, as well as allowing for faster iteration, is scaling down the current dataset sizes and models. For example, Kaddour (2023) demonstrate that it is possible to train a model on 745x less data, but lose only 1.9% performance on GLUE (Wang et al., 2018a) and 2.5% performance on super-natural instructions (Wang et al., 2022b). Additionally, Polo et al. (2024) show that a wide variety of evaluation benchmarks can be distilled into only 100 data points, while maintaining an accuracy within  2% of the true accuracy. One area that is still uncertain is whether the results from small models scale up to larger models, or if there is any way to predict the results on large models, based on experiments from small models. This is an active area of research.
@@
Preliminary works have been done in this direction, including some works on data attribution and valuation (Ghorbani & Zou, 2019; Ilyas et al., 2022), and data measurements (Mitchell et al., 2022). Works on data valuation and attribution (discussed in §9) attempt to better understand how each data point affects a model’s predictions, but may be used as inspiration for methods that directly evaluate the data itself. Additionally, there are a number of metrics that directly measure the intrinsic characteristics of data including distance, density, diversity, tendencies (mean, median, mode), and association (correlation, mutual information), but these measures have yet to be connected with a model’s performance on downstream data (Mitchell et al., 2022).
Developing metrics that directly evaluate data can significantly decrease the time spent on method development and increase the amount of exploratory work possible. Furthermore, removing the need to train models (which can be very costly) can allow for lower-resourced organizations and individuals to contribute to the field, improving diversity and inclusion in the process of large model development.
@@
Another high-impact direction for increasing the rate of research on data selection is the development of data-centric benchmarks and challenges. One hurdle of advancing data selection is that works are done across a wide variety of settings and cannot be directly compared. Specifically, much of data selection research using heuristic filters is performed as part of collecting a new dataset. This makes it impossible to compare selection methods with another dataset as they may contain different quantities of data, and sometimes even use varying model architectures and sizes. In other domains, such as vision and vision+language, challenges and benchmarks have been developed that provide well-founded evaluation settings that allows researchers from many institutions to participate in a combined effort to push the frontier of knowledge and capabilities in data selection (discussed in “Increasing the rate of data selection research” in §8). Creating similar challenges for language models can go a long way in encouraging progress.
@@ 11.3 Shifting compute time from model training to data processing
It is now well known that training data plays a very significant role in the performance of a model, therefore creating an increased desirability to spend more compute on data processing and selection (and possibly away from model training). To improve beyond what is possible only through heuristics, it may be necessary to utilize more expensive methods.
One example of how this may be accomplished is through the use of methods which have previously only been used in low-data settings (e.g., alignment) and applying them to settings with greater amounts of data (e.g., pretraining). For example, methods inspired by model-based filtering, or reward model-reweighting techniques can be used to improve the “quality” of pretraining data. Additionally, techniques for generating synthetic data can be used to create more diverse training data (Bradley et al., 2023) and methods previously used only for red-teaming (Samvelyan et al., 2024) can be used to generate training data for more robust models. As compute gets cheaper, it becomes more feasible to spend the cost of directly applying the existing methods. However, another avenue worth exploring is using cheap approximations for utility functions, such as computing similarity using bag-of-n-grams and hash functions rather than from model representations.
Another method for improving data that would require large amounts of compute is additive data selection. Currently, most data selection methods (especially for pretraining) remove undesirable content from data points. However, it is possible that in the future it will be beneficial to have additive data selection methods. For example, it would be possible to use a language model to infill information within a data point. One example would be to include explanations for some phrases: e.g., “the basketball player wasn’t able to move due to the double dribble rule so he passed the ball to his teammate” where the italicized text has been added by an infilling model.
@@ Appendix A Heuristic Filtering Details
  • If the line is a counter (e.g., “3 likes”)
@@
  • At the beginning of the line (e.g., “sign-in”)
  • At the end of the line (e.g., “Read more…”)
  • Anywhere in the line (e.g., “items in cart”)
@@
  • have an word from the “List of Dirty, Naughty, Obscene, or Otherwise Bad Words” [^36] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words